OpenStudy (mind2000):

Estimate the diameter of a circle with a circumference of 56 feet.

mathslover (mathslover):

Circumference of a circle = \(2\pi r\) where r is the radius of the circle. Since, 2*r = diameter of the circle So , Circumference of a circle = \((2r) \pi = \textbf{Diameter} \times \pi\) Since, circumference = 56 feet. Therefore, 56 = diameter * \(\pi\) Thus, diameter = \(\cfrac{56}{\pi}\)